# Charged current neutrino scattering from nucleons

### v1

@@ III.2 Inelastic scattering processes
The results are presented in Fig. 8 for $\frac{d\sigma}{dQ^{2}}$ vs $Q^{2}$ for the charged current neutrino and antineutrino induced one pion production off the nucleon target. These results are presented on the left panel for $\nu_{\mu}p\longrightarrow\mu^{-}p\pi^{+}$ (top panel), $\nu_{\mu}n\longrightarrow\mu^{-}n\pi^{+}$ (central panel), $\nu_{\mu}n\longrightarrow\mu^{-}n\pi^{0}$ (bottom panel), and on the right panel for $\bar{\nu}_{\mu}p\longrightarrow\mu^{+}n\pi^{-}$ (top panel), $\bar{\nu}_{\mu}n\longrightarrow\mu^{+}p\pi^{-}$ (central panel), $\bar{\nu}_{\mu}p\longrightarrow\mu^{+}n\pi^{0}$ processes. In the case of single pion production, we have taken the contribution from spin $\frac{1}{2}$ resonances like $P_{11}(1440)$, $S_{11}(1535)$, $S_{31}(1620)$, $S_{11}(1650)$, and spin $\frac{3}{2}$ resonances like $P_{33}(1232)$, $D_{13}(1520)$, $D_{33}(1700)$, $P_{13}(1720)$. In the vector sector, the helicity amplitudes for all these resonance excitations are given by the MAID parameterization  Tiator:2011pw . In the case of spin $\frac{1}{2}$ resonances, the s-channel and u-channel hadronic currents for the positive and negative parity resonances are discussed in detail in Ref. SajjadAthar:2022pjt . We have also considered the nonresonant background contribution SajjadAthar:2022pjt . The Full model corresponds to the results when Eq. (27) is used in the evaluation of the matrix element, while $P_{33}(1232)$ corresponds to the results when only $P_{33}(1232)$ dominance is assumed. The results are presented with a cut of 1.4 GeV and 2.0 GeV on the center of mass energy $W$, while calculating the differential scattering cross section. It may be observed from the figure that in the case of $p\pi^{+}$ and $n\pi^{-}$ channels, which are purely isospin $I=\frac{3}{2}$ states, the cross section is dominated mainly by $P_{33}(1232)$ resonance and the contribution from the nonresonant terms as well as from the higher resonances is almost negligible in the kinematic region considered in this work. The overall cross section decreases when a cut of $W<1.4$ GeV is applied and in the peak region of $Q^{2}$, the cross section is decreased by about $23\%~{}(27\%)$ as compared to the results obtained when a cut of 2 GeV on $W$ is applied, for $p\pi^{+}~{}(n\pi^{-})$ channels. In the case when a cut of $W<1.4$ GeV is applied, the results obtained using the full model and only $P_{33}(1232)$ are almost the same as there is no resonance other than $P_{33}(1232)$ in the range of $W\sim 1.4$ GeV contributing to the pion production channel. In the case of $n\pi^{+}$ and $p\pi^{-}$ channels, we find some contribution from the nonresonant terms as well as from the higher resonances, when a cut of $W<2$ GeV is applied. A cut of $W<1.4$ GeV reduces the cross section by about $75\%~{}(78\%)$ for the $n\pi^{+}~{}(p\pi^{-})$ channels as compared to the results obtained when a cut of 2 GeV is applied on the center of mass energy. It may be observed from the figure that the presence of the nonresonant terms and the higher resonances leads to an enhancement in the cross section, which is about 15–20% in the peak region of $Q^{2}$, in the case of $p\pi^{0}$ and $n\pi^{0}$ channels. Moreover, we observe a suppression in the cross section, which is about $50\%~{}(55\%)$ for the $p\pi^{0}~{}(n\pi^{0})$ channels when a cut of 1.4 GeV is applied instead of a cut of 2 GeV on the center of mass energy $W$.
@@
In Fig. 9, the results are presented for $\frac{d\sigma}{dQ^{2}}$ vs $Q^{2}$ for the charged current $\nu_{\mu}$ (left panel) and $\bar{\nu}_{\mu}$ (right panel) induced one kaon production off the nucleon target. These results are presented on the left panel for the processes $\nu_{\mu}N\longrightarrow\mu^{-}N^{\prime};K^{i}~{}~{}N,N^{\prime}=n$ or $p$; $K^{i}=K^{+}$ or $K^{0}$, where only the nonresonant background contributions have been considered as there is no established resonance which decays to a kaon and a nucleon in the region of $W<2$ GeV. On the right panel, the results are presented for the processes $\bar{\nu}_{\mu}N\longrightarrow\mu^{+}N^{\prime};K^{i}~{}~{}N,N^{\prime}=n$ or $p$; $K^{i}=K^{-}$ or $\bar{K}^{0}$, we have taken the contribution from the nonresonant terms as well as from the $S=-1$ $\Sigma^{\star}(1385)$ resonance. The results are presented with a cut of 1.6 GeV and 2.0 GeV on $W$, while calculating the differential scattering cross section. It may be observed from the figure that the effect of $W$ cut is quite important for the single kaon production processes. In the case of neutrino induced single kaon production, when the cut on $W$ is decreased from 2 GeV to 1.6 GeV, the cross section for all three processes decreases by about 75–80%, while in the case of antineutrino induced antikaon production processes, this reduction is about 60–65% for the three different reactions viz. $\bar{\nu}_{\mu}+p\longrightarrow\mu^{+}+p+K^{-}$, $\bar{\nu}_{\mu}+p\longrightarrow\mu^{+}+n+\bar{K}^{0}$ and $\bar{\nu}_{\mu}+n\longrightarrow\mu^{+}+n+K^{-}$.
Fig. 10 presents the results for $\frac{d\sigma}{dQ^{2}}$ vs $Q^{2}$ for the charged current $\nu_{\mu}$ (left panel) and $\bar{\nu}_{\mu}$ (right panel) induced single eta production off the nucleon target. These results are presented on the left panel for $\nu_{\mu}n\longrightarrow\mu^{-}p\eta$ and on the right panel for $\bar{\nu}_{\mu}p\longrightarrow\mu^{+}n\eta$. We have taken the contribution from the spin $\frac{1}{2}$ resonances like $S_{11}(1535)$, $S_{11}(1650)$, $P_{11}(1710)$, $P_{11}(1880)$, and $S_{11}(1895)$ along with the s-channel and u-channel Born terms. The $\eta$ production is dominated by the excitation of $S_{11}(1535)$ resonance and its subsequent decay to $\eta N$. To show the dominance of $S_{11}(1535)$ resonance in the case of $\eta$ production, we have also presented the results by taking into account the contribution from $S_{11}(1535)$ resonance only. The results are presented with a cut of 1.6 GeV and 2.0 GeV on $W$, while calculating the differential scattering cross section. From the figure, one may notice that in the case of $W<1.6$ GeV cut, the results obtained with the full model and $S_{11}(1535)$ resonance only overlap, as there is no other resonance contributing in this region of $W$ and the contribution of the nonresonant terms in the case of $\eta$ production is very small. As the cut on $W$ is increased from 1.6 GeV to 2 GeV, the differential cross section increases, and in the peak region of $Q^{2}$, this increase in the cross section is about 35–40% for both the neutrino and antineutrino induced $\eta$ production processes.
@@ III.3 Deep inelastic scattering
In Fig. 12, we have presented the results for $\frac{d\sigma}{dQ^{2}}$ vs $Q^{2}$ for the charged current $\nu_{\mu}$ (left panel) and $\bar{\nu}_{\mu}$ (right panel) induced deep inelastic scattering off the nucleon target at $E_{{\nu_{\mu}}{\bar{\nu}_{\mu}}}$=3 GeV. This is because considering $W>1.8$ or $2$ GeV at $E_{{\nu_{\mu}}{(\bar{\nu}_{\mu})}}$=2 GeV would result a small contribution in the lower region of $Q^{2}$ only. These numerical results are obtained by incorporating the higher order perturbative corrections up to next-to-the-leading order(NLO) as well as the nonperturbative effects like target mass correction and higher twist correction up to twist-4. These higher order perturbative and nonperturbative corrections are discussed in detail in Ref. Zaidi:2019asc . It may be noticed that the numerical results obtained with a lower cut on $W$, i.e., $W>1.8$ GeV span over the wider kinematic range of $Q^{2}$ as well as are higher than the results obtained with a cut of $W>2$ GeV. For example, the neutrino-nucleon (antineutrino-nucleon) differential cross section with a cut of $W>2$ GeV gets suppressed by 40% (44%) at $Q^{2}=1$ GeV2, 45% (50%) at $Q^{2}=1.5$ GeV2 and 60% (64%) at $Q^{2}=2$ GeV2 as compared to the results obtained with $W>1.8$ GeV cut. It may be observed that the effect of center of mass energy cut is more pronounced in the peak region as well as it excludes the region of high $x$ relevant to the tail of the differential cross sections, where the nonperturbative effects are significant. Furthermore, we find that the differential cross section gets reduced due to the inclusion of higher twist corrections in the entire range of $Q^{2}$. This reduction in the differential cross section is more pronounced for the lower values of $Q^{2}$ which gradually becomes small with the increase in $Q^{2}$, for example, there is a reduction of about 16% at $Q^{2}=1$ GeV2, 14% at $Q^{2}=1.5$ GeV2 and 12% at $Q^{2}=2$ GeV2 in the $\nu_{\mu}$ induced deep inelastic cross sections. It is important to point out that the inclusion of higher twist (HT) correction in $\bar{\nu}_{\mu}-N$ deep inelastic cross sections results in an enhancement, which is contrary to the behavior of HT correction observed in the case of $\nu_{\mu}-N$ DIS reaction. Quantitatively, this enhancement in the numerical results incorporating a cut of $W>2$ GeV is about 14% at $Q^{2}=1$ GeV2 which increases to 30% at $Q^{2}=1.5$ GeV2.
@@ IV Summary and conclusions
  • The results for the antineutrino cross sections are qualitatively similar but smaller as compared to the neutrino scattering cross sections. However, in this case the additional channel of hyperon production $\Lambda$, $\Sigma$ is opened which is very useful in doing SCC and T-invariance studies as the effects are larger and can be pursued by studying the angular distribution of pions coming from $\Lambda(\Sigma)\to N\pi$ decays.
@@
  • The effect of CM energy cut of $W>1.8$ GeV or $W>2$ GeV makes a significant difference on the cross sections. Therefore, it is very important that an appropriate cut on $W$ is imposed to identify the true DIS events in the (anti)neutrino-nucleon scattering.
The results discussed in this paper will be quite useful to estimate the $\nu_{l}(\bar{\nu}_{l})-N$ scattering cross sections, particularly for the experiments, which are studying $Q^{2}$-distribution in the few GeV energy region.

### v2

@@ III.2 Inelastic scattering processes
Figure 12: $\sigma$ vs. $E_{\nu_{\mu}}$ for the processes $\nu_{\mu}p\longrightarrow\mu^{-}p\pi^{+}$ (left panel) and $\nu_{\mu}n\longrightarrow\mu^{-}p\pi^{0}$ (right panel) with $W<1.4$ GeV. Solid line is the result of the present model with deuteron effect; compared with other theoretical models like DCC Nakamura:2015rta  (dashed line), HNV Hernandez:2007qq  (dashed-dotted line), Hybrid Gonzalez-Jimenez:2016qqq  (double-dotted-dashed line), NuWro Gonzalez-Jimenez:2016qqq  (double-dashed-dotted line) and Giessen Lalakulich:2010ss  (dotted line). Data points quoted as ANL extracted and BNL extracted are the reanalyzed data by Wilkinson et al. Wilkinson:2014yfa and Rodrigues et al. Rodrigues:2016xjj .
@@
The results are presented in Fig. 11 for $\frac{d\sigma}{dQ^{2}}$ vs $Q^{2}$ for the charged current neutrino and antineutrino induced one pion production off the nucleon target. These results are presented on the left panel for $\nu_{\mu}p\longrightarrow\mu^{-}p\pi^{+}$ (top panel), $\nu_{\mu}n\longrightarrow\mu^{-}n\pi^{+}$ (central panel), $\nu_{\mu}n\longrightarrow\mu^{-}n\pi^{0}$ (bottom panel), and on the right panel for $\bar{\nu}_{\mu}p\longrightarrow\mu^{+}n\pi^{-}$ (top panel), $\bar{\nu}_{\mu}n\longrightarrow\mu^{+}p\pi^{-}$ (central panel), $\bar{\nu}_{\mu}p\longrightarrow\mu^{+}n\pi^{0}$ processes. In the case of single pion production, we have taken the contribution from spin $\frac{1}{2}$ resonances like $P_{11}(1440)$, $S_{11}(1535)$, $S_{31}(1620)$, $S_{11}(1650)$, and spin $\frac{3}{2}$ resonances like $P_{33}(1232)$, $D_{13}(1520)$, $D_{33}(1700)$, $P_{13}(1720)$. In the vector sector, the helicity amplitudes for all these resonance excitations are given by the MAID parameterization  Tiator:2011pw . In the case of spin $\frac{1}{2}$ resonances, the s-channel and u-channel hadronic currents for the positive and negative parity resonances are discussed in detail in Ref. SajjadAthar:2022pjt . We have also considered the nonresonant background contribution SajjadAthar:2022pjt . The Full model corresponds to the results when Eq. (27) is used in the evaluation of the matrix element, while $P_{33}(1232)$ corresponds to the results when only $P_{33}(1232)$ dominance is assumed. The results are presented with a cut of 1.4 GeV and 2.0 GeV on the center of mass energy $W$, while calculating the differential scattering cross section. It may be observed from the figure that in the case of $p\pi^{+}$ and $n\pi^{-}$ channels, which are purely isospin $I=\frac{3}{2}$ states, the cross section is dominated mainly by $P_{33}(1232)$ resonance and the contribution from the nonresonant terms as well as from the higher resonances is almost negligible in the kinematic region considered in this work. The overall cross section decreases when a cut of $W<1.4$ GeV is applied and in the peak region of $Q^{2}$, the cross section is decreased by about $23\%~{}(27\%)$ as compared to the results obtained when a cut of 2 GeV on $W$ is applied, for $p\pi^{+}~{}(n\pi^{-})$ channels. In the case when a cut of $W<1.4$ GeV is applied, the results obtained using the full model vs. considering only $P_{33}(1232)$ contribution are almost the same as there is no resonance other than $P_{33}(1232)$ in the range of $W\sim 1.4$ GeV contributing to the pion production channel. In the case of $n\pi^{+}$ and $p\pi^{-}$ channels, we find some contribution from the nonresonant terms as well as from the higher resonances, when a cut of $W<2$ GeV is applied. A cut of $W<1.4$ GeV reduces the cross section by about $75\%~{}(78\%)$ for the $n\pi^{+}~{}(p\pi^{-})$ channels as compared to the results obtained when a cut of 2 GeV is applied on the center of mass energy. It may be observed from the figure that the presence of the nonresonant terms and the higher resonances leads to an enhancement in the cross section, which is about 15–20% in the peak region of $Q^{2}$, in the case of $p\pi^{0}$ and $n\pi^{0}$ channels. Moreover, we observe a suppression in the cross section, which is about $50\%~{}(55\%)$ for the $p\pi^{0}~{}(n\pi^{0})$ channels when a cut of 1.4 GeV is applied instead of a cut of 2 GeV on the center of mass energy $W$.
In Fig. 12, we have compared the theoretical results for the single pion production induced by neutrinos when a cut of 1.4 GeV is applied on $W$, obtained in the different models like the present model, the dynamical coupled channel (DCC) model by Nakamura et al. Nakamura:2015rta , the Hernandez, Nieves, and Valverde (HNV) model Hernandez:2007qq , the extension of HNV model by incorporating Regge model at high energies (Hybrid) by Gonzalez-Jimenez et al. Gonzalez-Jimenez:2016qqq , the results from NuWro Gonzalez-Jimenez:2016qqq Monte Carlo generator, and the Giessen model by Lalakulich et al. Lalakulich:2010ss . It may be observed from the figure that in the case of $p\pi^{+}$ channel, the results obtained in our model are quite consistent with the results obtained by the hybrid model and are in a very good agreement with the reanalyzed data of ANL Radecky:1981fn and BNL Kitagaki:1986ct by Wilkinson et al. Wilkinson:2014yfa and Rodrigues et al. Rodrigues:2016xjj . However, the results obtained in the other models like DCC, HNV, etc., are higher than the results obtained by us as well as the experimental data, but are consistent with one another. Moreover, the results obtained by the Giessen group Lalakulich:2010ss are lower than our results. In the case of $p\pi^{0}$ channel, our results are in a quite good agreement with the experimental data, while the results obtained in the other theoretical models are higher than our results. At energies $E_{\nu_{\mu}}<0.8$ GeV, the results obtained in the various models are consistent with each other. It may be noticed from the figure that there is a large difference among the various theoretical models and Monte Carlo generators available in the literature. In order to understand the dynamics of the single pion production, which is the simplest inelastic process, further theoretical and experimental work is required.
@@
In Fig. 13, the results are presented for $\frac{d\sigma}{dQ^{2}}$ vs $Q^{2}$ for the charged current $\nu_{\mu}$ (left panel) and $\bar{\nu}_{\mu}$ (right panel) induced one kaon production off the nucleon target. These results are presented on the left panel for the processes $\nu_{\mu}N\longrightarrow\mu^{-}N^{\prime}K^{i};~{}~{}N,N^{\prime}=n$ or $p$; $K^{i}=K^{+}$ or $K^{0}$, where only the nonresonant background contributions have been considered as there is no established resonance which decays to a kaon and a nucleon in the region of $W<2$ GeV. On the right panel, the results are presented for the processes $\bar{\nu}_{\mu}N\longrightarrow\mu^{+}N^{\prime}K^{i};~{}~{}N,N^{\prime}=n$ or $p$; $K^{i}=K^{-}$ or $\bar{K}^{0}$, we have taken the contribution from the nonresonant terms as well as from the $\Sigma^{\star}(1385)$ ($S=-1$) resonance. The results are presented with a cut of 1.6 GeV and 2.0 GeV on $W$, while calculating the differential scattering cross section. It may be observed from the figure that the effect of $W$ cut is quite important for the single kaon production processes. In the case of neutrino induced single kaon production, when the cut on $W$ is decreased from 2 GeV to 1.6 GeV, the cross section for all three processes decreases by about 75–80%, while in the case of antineutrino induced antikaon production processes, this reduction is about 60–65% for the three different reactions viz. $\bar{\nu}_{\mu}+p\longrightarrow\mu^{+}+p+K^{-}$, $\bar{\nu}_{\mu}+p\longrightarrow\mu^{+}+n+\bar{K}^{0}$ and $\bar{\nu}_{\mu}+n\longrightarrow\mu^{+}+n+K^{-}$.
Figure 16: Neutrino-nucleon structure functions $F_{2N}(x,Q^{2})$ (left panel) and $xF_{3N}(x,Q^{2})$ (right panel) vs. $x$ at $Q^{2}=2$ GeV2 and 10 GeV2. These results are obtained at NLO using different PDFs parameterizations Gluck:2007ck ; Nadolsky:2008zw ; Martin:2009iq ; Harland-Lang:2014zoa ; Dulat:2015mca .
[…]
@@ III.3 Deep inelastic scattering
Figure 18: $\frac{d\sigma}{dy}$ vs. $y$ for the deep inelastic scattering processes induced the neutrino (solid lines) and antineutrino (dashed lines) at $E_{\nu}=160$ GeV. These results are obtained using NLO, TMC, and HT effects, and are compared with the experimental data from CHARM collaboration CHARM:1981zpm .
In Fig. 16, we present the results for the free nucleon structure functions (Eq. 37), $F_{2N}(x,Q^{2})$ (left panel) and $xF_{3N}(x,Q^{2})$ (right panel) as functions of $x$, obtained using various PDF parameterizations, like GJR Gluck:2007ck , CTEQ6.6 Nadolsky:2008zw , MSTW Martin:2009iq , MMHT Harland-Lang:2014zoa and CT14 Dulat:2015mca at $Q^{2}=2$ GeV2 and 10 GeV2. This comparison highlights the dependence of nucleon structure functions on the choice of PDF parameterizations. The structure functions were evaluated at the next-to-leading order (NLO) following the methodologies prescribed by Vermaseren et al. Vermaseren:2005qc and Moch et al. Moch:2008fj . From the left panel, it is evident that the results obtained using MSTW, MMHT, and CT14 parameterizations are in close agreement across the entire range of $x$ for both $Q^{2}$ values. However, notable deviations are observed in the low to intermediate $x$ region ($x\leq$ 0.3) when comparing these results with those derived from the CTEQ6.6 Nadolsky:2008zw and GJR Gluck:2007ck parameterizations. These numerical results become consistent among themselves with the increase in $x$ and $Q^{2}$. From the right panel of the figure, it may be noticed that the results of $xF_{3N}(x,Q^{2})$ obtained by using GJR Gluck:2007ck and MMHT Harland-Lang:2014zoa parameterizations are significantly different from the results obtained by using CTEQ6.6 Nadolsky:2008zw , MSTW Martin:2009iq and CT14 Dulat:2015mca parameterizations in the region of low to intermediate $x$ and low $Q^{2}$. However, at $Q^{2}=10$ GeV2 this difference becomes small and the numerical results for different PDFs parameterizations are found to be in good agreement in the entire range of $x$, except for the results using GJR Gluck:2007ck parameterization in the region of low to intermediate $x$. This uncertainty arising due to the PDFs dependence of nucleon structure functions would further translate in the determination of neutrino-nucleon scattering cross sections in the kinematic region of low to intermediate $x$ and at the moderate values of $Q^{2}$. Hence, parton distribution functions are needed to be well determined in this kinematic region of $x$ and $Q^{2}$.
@@
In Fig. 17, we have presented the results for $\frac{d\sigma}{dQ^{2}}$ vs $Q^{2}$ for the charged current $\nu_{\mu}$ (left panel) and $\bar{\nu}_{\mu}$ (right panel) induced deep inelastic scattering off the nucleon target at $E_{{\nu_{\mu}}{\bar{\nu}_{\mu}}}$=3 GeV. This is because considering $W>1.8$ or $2$ GeV at $E_{{\nu_{\mu}}{(\bar{\nu}_{\mu})}}$=2 GeV would result a small contribution in the lower region of $Q^{2}$ only. These numerical results are obtained by incorporating the higher order perturbative corrections up to next-to-the-leading order (NLO) as well as the nonperturbative effects like target mass correction (TMC) and higher twist (HT) correction up to twist-4. These higher order perturbative and nonperturbative corrections are discussed in detail in Ref. Zaidi:2019asc ; Ansari:2020xne . It may be noticed that the numerical results obtained with a lower cut on $W$, i.e., $W>1.8$ GeV span over the wider kinematic range of $Q^{2}$ as well as are higher than the results obtained with a cut of $W>2$ GeV. For example, the neutrino-nucleon (antineutrino-nucleon) differential cross section with a cut of $W>2$ GeV gets suppressed by 40% (44%) at $Q^{2}=1$ GeV2, 45% (50%) at $Q^{2}=1.5$ GeV2 and 60% (64%) at $Q^{2}=2$ GeV2 as compared to the results obtained with $W>1.8$ GeV cut. It may be observed that the effect of center of mass energy cut is more pronounced in the peak region as well as it excludes the region of high $x$ relevant to the tail of the differential cross sections, where the nonperturbative effects are significant. Furthermore, we find that the differential cross section gets reduced due to the inclusion of higher twist corrections in the entire range of $Q^{2}$. This reduction in the differential cross section is more pronounced for the lower values of $Q^{2}$ which gradually becomes small with the increase in $Q^{2}$, for example, there is a reduction of about 16% at $Q^{2}=1$ GeV2, 14% at $Q^{2}=1.5$ GeV2 and 12% at $Q^{2}=2$ GeV2 in the $\nu_{\mu}$ induced deep inelastic cross sections. It is important to point out that the inclusion of higher twist (HT) correction in $\bar{\nu}_{\mu}-N$ deep inelastic cross sections results in an enhancement, which is contrary to the behavior of HT correction observed in the case of $\nu_{\mu}-N$ DIS reaction. Quantitatively, this enhancement in the numerical results incorporating a cut of $W>2$ GeV is about 14% at $Q^{2}=1$ GeV2 which increases to 30% at $Q^{2}=1.5$ GeV2.
In Fig. 18, the numerical results are presented for $\frac{1}{E_{\nu}}\frac{d\sigma}{dy}$ vs $y$ at $E_{\nu}=160$ GeV for the charged current $\nu_{\mu}-N$ (solid line) and $\bar{\nu}_{\mu}-N$ (dashed line) deep inelastic scattering cross sections. These numerical results are obtained at NLO incorporating the nonperturbative corrections like TMC following the operator product expansion approach and higher twist correction up to twist-4 following the renormalons approach Zaidi:2019asc ; Ansari:2020xne . We have found that at this (anti)neutrino energy these nonperturbative corrections are almost negligible (not shown here explicitly). The antineutrino-nucleon cross sections are lower compared to the neutrino-nucleon cross sections, with the degree of suppression increasing as $y$ increases. As observed from the figure, the numerical results align well with the CHARM experimental data CHARM:1981zpm , except in the peak region where slight deviations are evident.
@@ IV Summary and conclusions
  • The results for the antineutrino cross sections are qualitatively similar but smaller as compared to the neutrino scattering cross sections. However, in this case the additional channel of hyperon production $\Lambda$, $\Sigma$ is opened which is very useful in doing SCC and T-invariance studies as the effects are larger and can be pursued by studying the angular distribution of pions coming from $\Lambda(\Sigma)\to N\pi$ decays. Moreover, the study of SCC is quite relevant in a broader manner as the presence of SCC would directly imply beyond the standard model physics. The collider experiments like CMS CMS:2018wjk , Belle Belle:2018ttu , LHCb LHCb:2013hzx , etc. have reported the results for the transverse component of $\Lambda$, $\Lambda_{b}$, $\Lambda_{c}$ hyperons, which is the direct evidence of physics beyond the standard model. Therefore, our results on the weak production of $\Lambda$ hyperons and its polarization observables would provide an alternative method to study physics beyond the standard model.
@@
[…]
The results discussed in this paper will be quite useful to estimate the $\nu_{l}(\bar{\nu}_{l})-N$ scattering cross sections, particularly for the experiments, which are studying $Q^{2}$-distribution in the few GeV energy region. The MicroBooNE collaboration has reported the cross section results for the single hyperon production and $\eta$ meson production processes and plans to study the single kaon production and associated particle production processes in the near future. Moreover, the DUNE experiment also plans to perform cross section measurements in the neutrino energy region of about 2 GeV. Therefore, the results discussed in this paper would be useful in the analysis of the accelerator neutrino experiments like MicroBooNE, NOvA, DUNE, etc., as well as the other experiments with atmospheric neutrinos.